OpenStudy (amistre64):

How would you find the 1/3 derivative of: y=x^2 ?

OpenStudy (anonymous):

As in 1/3 is the constant that is outside the derivative?

OpenStudy (saifoo.khan):

\[\Large \frac{2x}{\frac13}\]?something like that?

OpenStudy (amistre64):

\[\frac{d^{1/3}}{dx^{1/3}}x^2\]

OpenStudy (anonymous):

saifoo.khan is right

OpenStudy (anonymous):

Look at section: Fractional derivative of a simple function

OpenStudy (amistre64):

Yeah, i knew it had something to do with the Gamma function :)

OpenStudy (saifoo.khan):

FORREAL? @myko

OpenStudy (amistre64):

so how would we go about implementing that?

OpenStudy (saifoo.khan):

No idea.

OpenStudy (amistre64):

\[D^{1/3}[x^2]=\frac{\Gamma(2+1) }{\Gamma(2-\frac{1}{3}+1)}x^{2-\frac{1}{3}}\]

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=gamma%282%29%2Fgamma%283-1%2F3%29+x%5E%7B2-1%2F3%7D i get this; but i still cant figure out how its spose to work